A man received a 9 months loan 5000$ from a bank of the proceed were 4500 what is discount rate


1
Expert's answer
2021-06-16T15:17:06-0400

Discount rate =OriginalamountActualamountOriginalamount\frac {Original amount - Actual amount} {Original amount}


Discount Rate =500045005000×100=10%\frac {5000-4500}{5000}×100 =10\%